Solve the system by using the addition method. -0.1x+0.05y=5 and 0.05x+0.06y=2.7 What is x?
Answer by funmath(2933)

Solve the system by using the addition method.
L1) -0.1x+0.05y=5
L2) 0.05x+0.06y=2.7
:
It's probably easiest to eliminate all of the decimals and solve as usual. Multiply L1 and L2 by 100
100(-0.1x+0.05y=5)-------->-10x+5y=500
100(0.05x+0.06y=2.7)------>5x+6y=270
Now multiply your new L2 by 2 and add the two equations together:
L2) 2(5x+6y=270)----->10x+12y=540
:
-10x+5y=500
+10x+12y=540
________________
0x+17y=1040
17y=1040
17y/17=1040/17
y=61.17647059 Round off to whatever your teacher asks for.
Plug that in for y in one of the equations and solve for x.
-0.1x+0.05(61.17647059)=500
-0.1x+3.0588235=5
-0.1x=5-3.0588235
-0.1x=1.9411765
-0.1x/-0.1=1.9411765/-0.1
x=-19.411765
The solultion is (-19.411765,61.176470), if I was your teacher I'd probably have you round it off to: (-19.41,61.18)
Happy Cacluating!!!